How do you solve (sin120+2cos120)/ sec300 without using a calculator and using special values?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52879)

.

From your post, it is unclear to me, if using special values is allowed.



Answer by Alan3354(69443)

How do you solve (sin120+2cos120)/ sec300 without using a calculator and using special values?
========================
There is nothing to be solved - no variable, no unknowns.
----
It can be evaluated.
(sin(120) + 2cos(120))/sec(300) = (sin(120) + 2cos(120))*cos(300)
= ((sqrt(3)/2) + 2*(-0.5))*0.5
= (sqrt(3)/2) - 1)*0.5
= sqrt(3)/4 - 0.5